OpenStudy (anonymous):

Given right triangle ABC. In the given right triangle, find the length of the hypotenuse, AB, if BC = 6 and AC = 5. sqrt 36 sqrt 61 sqrt 11

OpenStudy (sleepyjess):

Do you know the pythagorean theorem?

OpenStudy (anonymous):

a^2 + b^2 = c^2

OpenStudy (anonymous):

6^2 + 5^2 = c^2

OpenStudy (anonymous):

take it from there M'lady

OpenStudy (anonymous):

so sqrt 11?

OpenStudy (sleepyjess):

No, what is \( 6^2\)?

OpenStudy (anonymous):

wait i always think its plus lol

OpenStudy (anonymous):

to the power it means the number times itself

OpenStudy (anonymous):

so 6^2 means 6 * 6

OpenStudy (sleepyjess):

Here is the pythagorean theorem: \(\sf a^2+b^2=c^2\). Substitute in 6 for a, and 5 for b, then solve.

OpenStudy (sleepyjess):

Ok, so you are confused on exponents? x^2 means x*x x^3 means x*x*x etc

OpenStudy (anonymous):

In many cases jess you won't be able to display a perfect perfect power so it's important that she realizes ^ is "to the power of"

OpenStudy (sleepyjess):

What do you mean a perfect perfect power?

OpenStudy (anonymous):

ok got that point lol thnx

OpenStudy (anonymous):

I meant to only put one perfect my bad.

OpenStudy (sleepyjess):

Ok, so what do you get once you add 6^2 + 5^2?

OpenStudy (anonymous):

It means that you won't always be able to make a fancy version of 6^2 and so she'll need to know what ^ means

OpenStudy (sleepyjess):

Oh, like this: \(\sf 6^2\) vs this? 6^2

OpenStudy (anonymous):

yea

OpenStudy (sleepyjess):

ok :)

OpenStudy (anonymous):

6^2 = 36 5^2 = 25 36 + 25 = ?

OpenStudy (anonymous):

equalllllll....????????

OpenStudy (anonymous):

Basic math hereeeee

OpenStudy (anonymous):

61

OpenStudy (anonymous):

there you go

OpenStudy (anonymous):

now

OpenStudy (anonymous):

square root of 61 isssss?

OpenStudy (sleepyjess):

Put into a calculator \(\sqrt{61}\)

OpenStudy (anonymous):

Here you can use this M'lady http://www.wolframalpha.com/input/?i=square+root+61

OpenStudy (sleepyjess):

That works too

OpenStudy (anonymous):

as a calculator I mean

OpenStudy (anonymous):

thnx

OpenStudy (sleepyjess):

So, in the end we have \(\sf\sqrt{61}\approx 7.8\)

OpenStudy (anonymous):

tada~